### v1

# Comparing Hallucination Detection Metrics for Multilingual Generation

Abstract
While many automatic hallucination detection techniques have been proposed for English texts, their effectiveness in multilingual contexts remains unexplored. This paper aims to bridge the gap in understanding how these hallucination detection metrics perform on non-English languages. We evaluate the efficacy of various detection metrics, including lexical metrics like ROUGE and Named Entity Overlap and Natural Language Inference (NLI)-based metrics, at detecting hallucinations in biographical summaries in many languages; we also evaluate how correlated these different metrics are to gauge whether they measure the same phenomena. Our empirical analysis reveals that while lexical metrics show limited effectiveness, NLI-based metrics perform well in high-resource languages at the sentence level. In contrast, NLI-based metrics often fail to detect atomic fact hallucinations. Our findings highlight existing gaps in multilingual hallucination detection and motivate future research to develop more robust detection methods for LLM hallucination in other languages.
1 Introduction
Large Language Models (LLMs) have brought about remarkable advances in text generation. However, they are still prone to factuality hallucination - the generated text conflicts with established world knowledge Huang et al. (2023); Zhang et al. (2023). Despite considerable research efforts towards understanding and detecting hallucinations, the focus has predominantly been on English texts Huang et al. (2023); Zhang et al. (2023); Ji et al. (2023). This emphasis has resulted in a significant knowledge gap regarding hallucinations in multilingual contexts, and it is currently unclear whether the methods developed for detecting and addressing hallucinations in English are effective or even applicable in multilingual settings.
In this paper, we address the highlighted issues by evaluating the effectiveness of various metrics (initially proposed for English generation) within a multilingual context. Our approach involves a comparative analysis of traditional lexical metrics, such as ROUGE Lin (2004) and Named Entity Overlap, alongside Natural Language Inference (NLI) metrics. Additionally, we compare reference-based metrics with pairwise metrics that are based on the consistency among generated samples; we also present a correlation study between these automated metrics and human factuality verification. These evaluations shed light on the efficacy of these metrics in non-English languages and aim to establish a more robust framework for hallucination detection in multilingual language models.
We empirically evaluate different hallucination detection techniques in the multilingual context and find that:
  • Lexical overlap metrics (e.g., ROUGE, named entity overlap) do not correlate with NLI-based metrics in detecting factual hallucinations in both reference-free and reference-based settings.
  • Pairwise NLI-based metrics strongly correlate with reference-based metrics in high-resource languages, but this significantly diminishes in low-resource languages.
  • Automatic NLI-based metrics effectively identify sentence-level hallucinations in high-resource languages when compared to human evaluations. However, their performance diminishes when assessing simpler atomic facts.
These findings indicate several open problems for multilingual hallucination detection. While traditional lexical overlap methods and pairwise comparisons of multiple generated texts are more accessible approaches for low-resource languages, they are often inadequate for hallucination detection. This highlights that the effectiveness of hallucination detection is closely tied to the availability and quality of language resources, following the trend observed in English that detection accuracy depends on natural language understanding abilities Manakul et al. (2023); Min et al. (2023). Our work points to a substantial gap in hallucination detection in multilingual contexts, necessitating focused future research to bridge this divide.
2 Analyzing Methods for Multilingual Hallucination Detection
Hallucination refers to machine-generated content that is not faithful to a reference. This reference could be source text, preceding generated context, or world knowledge Ji et al. (2023); Zhang et al. (2023). Hallucinations can be further classified into two types: verifiable hallucinations, which happen when the generated content directly contradicts the reference, and unverifiable hallucinations, meanwhile, refer to generated content that cannot be verified with the source material.
In this paper, we measure the efficacy of different metrics on detecting multilingual hallucinations. We focus on biography generation, a domain that is particularly sensitive to factual accuracy and coherence Min et al. (2023); Dhuliawala et al. (2023). We test a suite of automatic metrics, each of which caters to a different aspect of factual generation: ROUGE Lin (2004), named entity overlap, and Natural Language Inference (NLI)-based methods.
2.1 Multilingual Biography Generation
Table 1: Generation quality statistics for BLOOMZ-mt. The languages in bold are the most frequent in the ROOTS pretraining corpus Laurençon et al. (2022), and the underlined languages are covered in the xP3mt fine-tuning dataset Muennighoff et al. (2023). ”FLang.” refers to the most frequently generated language for each prompt language.
Lang. #Token #Sent. Valid % FLang. Acc.
en 78.3 2.64 99.97 en 96.0
zh 115.8 4.30 100.00 zh 92.43
es 62.8 2.01 100.00 es 92.33
fr 71.3 2.24 100.00 fr 93.23
vi 45.6 1.66 98.92 vi 71.67
id 46.3 1.76 98.30 en 36.45
de 63.3 2.33 99.58 en 2.79
it 58.1 1.94 99.76 en 3.31
ja 50.3 1.97 90.73 zh 21.85
bg 17.4 1.15 86.74 en 13.69
ro 9.6 0.93 80.24 en 2.68
sv 7.6 0.51 40.73 en 1.79
th 14.8 0.81 77.08 th 94.96
ru 10.2 0.68 55.49 ru 50.44
uk 5.7 0.40 35.24 uk 41.87
fa 3.2 0.13 10.80 ur 29.90
fi 1.7 0.11 9.76 fi 34.52
ko 2.0 0.09 8.37 ko 47.30
hu 0.8 0.05 6.28 pt 14.36
Avg. 34.8 1.35 31.65 - 50.01
Inspired by prior work measuring factuality in English Min et al. (2023), we generate parallel biographies in different languages. The generated texts are then compared against a reference text (for reference-based metrics) and other generated samples (pairwise metrics) to detect hallucinations.
This section characterizes the generation quality of these biographies (Table 1). We consider the average length of each biography (in tokens and sentences), along with estimates of how accurate the generation language is to the prompt language, as in some cases, multilingual LMs will generate continuations in an unexpected language Kang et al. (2023); Bawden and Yvon (2023).
The length of the generated texts varies notably across languages. High-resource languages like English and French average 78.3 and 115.8 tokens per response, respectively. However, mid-resource languages such as Thai tend to generate much shorter biographies and incomplete sentences, and low-resource languages fare even worse (for instance, Ukrainian averages just 5.7 tokens and 0.40 sentences), demonstrating the significant gap in generation abilities across languages.
We assess the accuracy of the generated languages through three metrics: the percentage of valid generations that is detectable for the langdetect package (Valid %) [^1] Some generations are incomplete, or their language is undetectable., the most frequently generated language for a given target language (Flang), and the accuracy of generated language out of the valid generations (Acc.). For high-resource languages like English, Chinese, Spanish, and French, the models generally generate text in the correct language; however, for the languages highlighted with an underwave the model generates in the wrong language the majority of the time. Often, this is due to the model generating in a closely related high-resource language. For languages such as Italian and Bulgarian, many inaccurate generations are in English. Similarly, Japanese generations often switch to Chinese when mistakes occur. Languages with more distinctive linguistic features —such as Thai’s unique script—facilitate more accurate model generations.
2.2 Automatic Methods for Multilingual Factuality Detection
Following the language verification and ruling out examples where the generations are in the wrong languages, the next phase involves the detection of hallucinations in a multilingual environment by assessing the consistency between a target generation and either a reference text or its other generations. This section considers various automatic metrics for detecting hallucinations in long-form generations.
ROUGE
The ROUGE metric is employed to assess the token-level similarity between texts. We consider the ROUGE 1 (R1), 2 (R2), L (RL), and Lsum (RLsum) scores of the generated text against the reference.
Named Entity Overlap (NEO)
We hypothesize that the sets of named entities in the gold and generated text will differ if there is hallucination in the generation Nan et al. (2021). We calculate the F1, precision, and recall scores of named entities between the generated and reference text as an estimate for factual hallucinations.
NLI-based Detection
Following Manakul et al. (2023) and Elaraby et al. (2023), we adopt the NLI-based zero-shot sentence-level SUMMAC ($SummaC_{zs}$) scoring system Laban et al. (2021) to evaluate hallucinations. The $SummaC_{zs}$ method was originally developed to gauge the consistency between a summary $S$ and a document $D$, by segmenting them into sentences $S_{1},\dots,S_{N}$ and $D_{1},\dots,D_{M}$ respectively. Aligning with the optimal configuration in Laban et al. (2021), we employ the $max$ operator to compute the score for a sentence. Denote $e_{S_{n}}^{D_{m}}$ and $c_{S_{n}}^{D_{m}}$ as the entailment and contradiction score for the generated sentence $S_{n}$ given the reference sentence $D_{m}$, respectively.
We define three metrics to quantify verifiable hallucination and one metric to quantify unverifiable hallucination, respectively. At sentence-level detection, for a generated sentence $S_{i}$ and a reference $D$, to detect verifiable hallucination, we define the following three metrics: $\mathbf{ENT_{S_{i}}}=\max_{m}{e_{S_{i}}^{D_{m}}}$, $\mathbf{CON_{S_{i}}}=\max_{m}{c_{S_{i}}^{D_{m}}}$, and $\mathbf{DIFF_{S_{i}}}=\max_{m}{e_{S_{i}}^{D_{m}}}-\max_{m}{c_{S_{i}}^{D_{m}}}$. To detect unverifiable hallucination, we define the following metric:
$\mathbf{UNV_{S_{i}}}=1-\max(\max_{m}{e_{S_{i}}^{D_{m}}},\max_{m}{c_{S_{i}}^{D_{m}}})$
When evaluating each of the above hallucination metrics on a generated text $\hat{t}$, we consider two settings as the reference text $t$:
Reference-based This setting compares $\hat{t}$ against the relevant biographical article in Wikipedia.
Pairwise We generate $k$ samples for each biography. In this setting, we compare $\hat{t}$ against the other generated samples for the same person and calculate the average score across all generations.
3 Experiment Setup
3.1 Dataset
Our curated dataset encompasses 19 languages: English, Spanish, Russian, Indonesian, Vietnamese, Persian, Ukrainian, Swedish, Thai, Japanese, German, Romanian, Hungarian, Bulgarian, French, Finnish, Korean, Italian, and Chinese. Using WikiData, we extract 500 human names that are covered by all of these languages on Wikipedia, based on diverse page view counts from 2022-01-01 to 2023-01-01. For our reference text, we use the Wikipedia API to obtain the full-page content. We detect instances where the LLMs generate text in an incorrect language with langdetect. [^2] APIs: https://query.wikidata.org/, https://pypi.org/project/wikipedia/, and https://pypi.org/project/langdetect/, respectively
3.2 Models and Prompting Details
In our experiments, we deploy the pretrained multilingual BLOOM model as well as the BLOOMZ-mt model, which is fine-tuned with machine-translated prompts Workshop (2023). We use nucleus decoding Holtzman et al. (2020) with $top\_p=0.9$ and generate five responses per prompt. For each evaluation language, we generate a prompt template with Google Translate. The template in English is ”Tell me a biography of ¡Name¿.”; all templates are in Appendix (Figure 2).
3.3 Detection Metric Details
ROUGE scores are calculated with TorchMetrics [^3] https://github.com/Lightning-AI/torchmetrics, and we remove all stopwords before calculating ROUGE-1. Entities are extracted with Spacy’s named entity recognizer [^4] https://spacy.io/api/entityrecognizer; we note that this tagger only covers 13 of the 19 languages considered in our experiments. For the NLI-based metric, we finetune the XLMR-large model Conneau et al. (2020) on the subset of the XNLI dataset Conneau et al. (2018) that intersects with the languages used in our experiments. The finetuned model has an average validation accuracy of 85.4% for the nine intersecting languages.
Language $\mathrm{ENT}$ $\mathrm{DIFF}$ $\mathrm{UNV}$
# examples in correct language ¿ 1,000
English 0.55 0.38 0.19
French 0.52 0.40 0.15
Chinese 0.56 0.41 0.21
Spanish 0.46 0.41 0.17
Thai 0.36 0.39 0.32
Vietnamese 0.35 0.31 0.00
Indonesian 0.28 0.31 0.09
# examples in correct language ¡ 1,000
Russian 0.16 0.21 0.11
Japanese 0.37 0.40 0.07
Ukrainian 0.23 0.19 0.17
Bulgarian 0.42 0.32 0.28
Korean 0.05 0.08 -0.01
# examples in correct language ¡ 100
Finnish 0.09 0.12 0.01
Italian 0.12 0.13 0.14
Persian 0.13 0.15 0.02
German 0.50 0.45 0.11
Romanian 0.00 0.00 0.14
Hungarian 0.24 0.21 0.11
Swedish 0.30 0.27 -0.29
Table 2: The correlation between the reference-based NLI result and the pairwise NLI result across different languages. The languages with underline are covered in the XNLI finetuing dataset. The numbers in yellow have the p-values larger than 0.05.
4 Results
In this section, we compare evaluation metrics to estimate hallucinations in our generated biographical corpus (Section 4.1). We then perform a correlation study to test whether the proposed metrics agree when hallucination occurs (Section 4.2). We find that traditional, reference-based lexical metrics (i.e., ROGUE and named entity overlap) are likely insufficient to capture hallucinations, as they rarely correlate with stronger NLI-based approaches. We also observe that pairwise metrics detect hallucinations less often in lower-resource languages when compared to reference-based metrics.
4.1 Comparison of Automatic Hallucination Metrics
Table 3: Results of different reference-based metrics for the BLOOMZ-mt model. ”-” indicates the language is not covered by the Spacy NER tool. All of the ROUGE and Named Entity Overlap (NEO) results are in percentage (%).
Language R1-F1 R1-P R1-R R2-F1 R2-P R2-R NEO-F1 NEO-P NEO-R $\mathrm{DIFF}$ $\mathrm{UNV}$
High-Resource Languages
English 1.83 87.58 0.94 0.87 47.38 0.44 4.27 53.41 2.26 -0.60 0.19
Chinese 6.43 57.34 3.76 2.07 23.22 1.17 4.69 35.27 2.79 -0.62 0.21
Spanish 2.77 85.86 1.47 1.35 49.10 0.71 3.28 48.48 1.76 -0.51 0.18
French 2.18 87.78 1.13 1.06 51.41 0.55 4.35 57.41 2.31 -0.54 0.16
Vietnamese 6.82 92.92 4.22 4.10 73.28 2.43 - - - -0.49 0.15
Indonesian 7.51 68.51 4.87 2.36 26.39 1.53 - - - -0.45 0.22
Middle-Resource Languages
German 0.38 71.34 0.19 0.11 35.22 0.05 0.83 36.06 0.42 -0.65 0.15
Italian 0.50 69.13 0.25 0.13 29.82 0.07 1.00 30.26 0.52 -0.58 0.17
Japanese 0.73 14.62 0.40 0.05 1.50 0.03 0.47 15.52 0.25 -0.72 0.21
Bulgarian 0.16 4.92 0.09 0.01 0.37 0.01 - - - -0.61 0.19
Romanian 1.02 69.75 0.53 0.42 45.79 0.22 0.39 17.47 0.20 -0.29 0.24
Swedish 0.66 86.37 0.33 0.32 77.23 0.16 1.28 45.24 0.66 -0.40 0.63
Low-Resource Languages
Thai 0.04 1.14 0.02 0.00 0.00 0.00 - - - -0.56 0.38
Russian 0.09 4.69 0.05 0.01 0.28 0.00 0.48 11.28 0.25 -0.58 0.47
Ukrainian 0.04 1.53 0.02 0.00 0.00 0.00 0.70 20.64 0.36 -0.53 0.66
Persian 0.00 0.00 0.00 0.00 0.00 0.00 - - - -0.50 0.92
Finnish 0.89 37.70 0.46 0.20 10.03 0.10 0.58 23.71 0.30 -0.59 0.91
Korean 0.18 6.58 0.09 0.01 0.88 0.00 0.24 8.48 0.12 -0.53 0.94
Hungarian 0.74 64.74 0.37 0.16 23.23 0.08 - - - -0.53 0.97
We first consider how different automatic methods for detecting hallucination perform across languages on the generated biographical data from the BLOOMZ-mt model (Table 3). High-resource languages, such as English, Chinese, Spanish, French, Vietnamese, and Indonesian, exhibit higher recall scores, which suggests that the text generated in these languages has better coverage of the corresponding Wikipedia reference content. This contrasts sharply with lower-resource languages, which demonstrate significantly diminished recall. Interestingly, languages that frequently produce incorrect language outputs (e.g., German and Italian) or often yield empty or incomplete generations (e.g., Swedish and Hungarian) still have relatively high precision scores. While these generations seem to contain few explicit hallucinations, they also exclude many facts from the reference, as indicated by their correspondingly low recall scores.
When considering the $\mathrm{DIFF}$ metrics, all languages under scrutiny resulted in negative scores, which was consistent even among the higher-resource languages like English and Chinese. This indicates a tendency towards contradictions in the generated text with their respective reference texts. For the $\mathrm{UNV}$ scores, higher and middle-resource languages (ranging from English to Romanian in the table 3) cluster within a similar scope of 0.15 to 0.25. In contrast, low-resource languages that often produce empty or incomplete generations, such as Ukrainian, Persian, Finnish, and Korean, show much higher $\mathrm{UNV}$ scores. This implies that the $\mathrm{UNV}$ metric is sensitive to incomplete text generations and missing information and may indicate the model’s generation errors beyond hallucination. As for the reference-free pairwise metrics, we observe similar trends across different languages (Table 9).
4.2 Correlation Study Across Metrics
In this section, we conduct a correlation analysis to determine how various metrics align in measuring hallucination in multilingual contexts. This includes (1) the correlation between lexical hallucination metrics and NLI-based metrics, (2) the performance of the four reference-based NLI metrics, and (3) the relationship between pairwise metrics and reference-based metrics.
Figure 1: Heat map of the Pearson Correlation between reference-based metrics averaged over high-resource languages. All the P-values are less than 0.05.
Entity Generation
Example of Annotation
Alessandro Del Piero
Gen: Alessandro Del Piero, born on September 28, 1976 in Brescia, Italy, is a former Italian professional football player who served as a forward. Wiki: Alessandro Del Piero, Italian male football player…The old Maldini, who was the head coach of the national team at the time, appointed newcomers Del Piero and Vieri as the main forwards… Comment: There are 4 facts in this sentence, with 1 contradictory hallucination and 1 unverifiable hallucination. The birth date is wrong; Wikipedia doesn’t mention the birth place; for the last entity, the evidence indirectly support it.
Example of Instruction-Conflict Hallucination
Carl Edward Sagan
Gen: Stephen Hawking (January 8, 1942 -) is a famous British physicist and cosmologist.
Example of Preceding-Context-Conflict Hallucination
Blanca of Castile
Gen: Blanca of Castile (Blanca of Castile), born in 1188 and died in 1252, was the princess of Castile, queen and regent of France. Blanca is a city in the Autonomous Community of Castilla-La Mancha, Spain.
Table 4: Examples of annotation process and types of hallucinations. Sentences highlighted in red represent verifiable hallucinations that contradict evidence found in Wikipedia; those marked in yellow denote unverifiable hallucinations which lack clear evidence for verification and content highlighted in green is considered factual.
Lexical hallucination metrics do not correlate with NLI-based metrics.
Figure 1 shows that in high-resource languages (i.e., English, Chinese, French, Spanish, Vietnamese, and Indonesian), ROUGE-1 and ROUGE-L metrics demonstrate a high degree of correlation, as do the ROUGE metrics and Named Entity Overlap (NEO). However, we generally find no correlation between lexical- and NLI-based metrics, indicating that while both lexical- and NLI-based approaches are commonly proposed as automatic methods for hallucination detection, they do not measure the same deviations from a reference text.
Reference-based NLI-based metrics.
We also observe interesting trends regarding the relationship between different NLI-based metrics (Figure 1). We find that $\mathrm{ENT}$ scores are highly correlated with the $\mathrm{DIFF}$ score, indicating that these metrics identify similar artifacts in the text. Surprisingly, no correlation is observed between $\mathrm{UNV}$ and $\mathrm{CON}$ scores. Moreover, we find a negative correlation between $\mathrm{UNV}$ and $\mathrm{CON}$ scores. This is because sentences that include verifiable hallucinations likely contradict the reference text. In contrast, sentences with information that is unsubstantiated by the reference (e.g., unverifiable) will be identified as neutral instead.
Pairwise and reference-based metrics correlate in higher-resource languages.
For high-resource languages in the XNLI finetuning dataset (English, French, Chinese, Spanish, Bulgarian, and German), we observe a higher correlation in the range of 0.35 to 0.56 for pairwise NLI metrics when it comes to detecting verifiable hallucinations by $\mathrm{ENT}$ score (Table 2). This suggests pairwise metrics can identify generated content that deviates from the reference. However, the Pearson Correlation Coefficient shows lower correlation values (in the range of 0.15 to 0.21) compared to pairwise $\mathrm{UNV}$. For lower-resource languages, such as Finnish, Italian, Persian, and others, the correlation coefficients with the entailment score are often in the range of 0.00 to 0.30 and not statistically significant, highlighting the challenge of effective hallucination detection in contexts with limited language resources.
5 Human Evaluation of Multilingual Hallucination Metrics
Currently, there are no standard multilingual hallucination detection datasets available on which to compare the considered methods. We, therefore, annotate model generations in English and Chinese; annotations are performed by native speakers of each language. They manually find all verifiable and unverifiable hallucinations by checking if the generation is supported by the Wikipedia reference at both the sentence- and atomic-fact-level. Examples of resulting annotations are shown in Table 4.
l—cccc—cccc
Sentence-Level Atomic-Fact-Level
[…]
Random - 10.84 82.86 50.00 - 52.1143.56 50.00
[…]
R2-P. 0.08 19.78 81.48 - 0.10 51.23 44.23 -
RL-P. 0.11 20.04 83.10 - 0.12 52.18 42.83 -
NEO-P. 0.14 17.49 80.84 - 0.09 53.09 45.32-
$\mathrm{DIFF}$ 0.21 38.46 89.49 68.21 0.19 57.46 54.41 56.26
$\mathrm{ENT}$ 0.31 40.32 90.86 70.97 0.23 60.71 57.48 53.47
$\mathrm{CON}$ 0.11 16.47 80.41 49.13 -0.01 51.49 52.16 50.48
Reference
R2-P. 0.21 30.05 89.08 - 0.19 53.28 46.25 -
RL-P. 0.17 28.54 85.35 - 0.13 50.31 49.93 -
NEO-P. 0.17 16.15 83.75 - 0.12 57.54 47.51 -
$\mathrm{DIFF}$ 0.34 56.11 94.14 75.00 0.31 65.85 60.90 63.19
$\mathrm{ENT}$ 0.49 65.32 94.96 78.97 0.35 68.00 63.6964.18
$\mathrm{CON}$ 0.08 31.56 87.49 69.77 -0.19 53.18 57.43 59.70
Table 5: Sentence- and atomic-fact-level verifiable hallucination detection in human evaluation. F denotes factual examples and NF denotes non-factual examples. The numbers in yellow have p-values of the correlation is larger than 0.05. Note that the correlation coefficient of $\mathrm{CON}$ should be expected as negative.